# Sergei Gukov, Ludmil Katzarkov, Josef Svoboda

### v2

@@ 1.2. Spectra in algebra and geometry
Since one of the predictions of the so-called 3d-3d correspondence is that a closed 3-manifold $Y$ corresponds to a VOA [8, 9, 69], for which $\widehat{Z}_{\sigma}$ is a character, it is natural to ask how the spectrum of conformal weights $\Delta_{\sigma}$ relates to other spectra mentioned above, defined more geometrically. Contrary to what one might naively expect, we find that the spectrum of conformal weights $\Delta_{\sigma}$ is rather different from the spectrum of the corresponding hypersurface singularity: while the former determines the leading $q$-power in the $q$-expansion, the latter encodes the structure of the $q$-series coefficients at large $q$-powers. In turn, the latter determines the behavior of the $q$-series near $q=e^{\hbar}\approx 1$ or, equivalently, the expansion in $\hbar$ near $\hbar=0$.
To summarize, expansion in $q$, as in $\widehat{Z}_{\sigma}(q)$, has direct contact with counting problems (curve counting, BPS states, etc.), with vertex algebras, and with the spectrum of conformal dimensions. These connections become less natural near $q=e^{\hbar}\approx 1$, where connections to complex Chern–Simons theory and the spectrum of hypersurface singularities become manifest. The two expansion limits are related by resurgent analysis, a powerful technique that, roughly speaking, allows to transfer (enlarge) the domain of a given function (or a power series, possibly, with zero radius of convergence).
This interplay between expansions near $q=0$ and $q=e^{\hbar}\approx 1$ plays an important role in curve counting [17, 28, 43, 38] and — since $\widehat{Z}_{\sigma}(q)$ admit an interpretation via curve counting too [35, 34, 21, 20] — in the study of $\widehat{Z}_{\sigma}(q)$-invariants [31, 3, 11, 13, 71]. In particular, near $q=e^{\hbar}\approx 1$ more natural objects are $Z_{a}(q)$, without a ‘hat’, and trans-series $Z_{\alpha}(q)$ of the complex Chern–Simons theory. This is consistent with what we find here: $Z_{0}$ and, more generally, $Z_{\alpha}$ appear to be more natural from the perspective of the singularity theory. And, perhaps not surprisingly (given the above explanations), the spectrum of the hypersurface singularity associated with $Y$ is closely related to the set of values of the classical Chern–Simons functional on $Y$, i.e., to the singularities on the Borel plane.
Acknowledgements. We would like to thank Denis Auroux, Martin Čech, Shimal Harichurn, Mrunmay Jagadale, Maxim Kontsevich, Slava Krushkal, Louisa Liles, András Némethi, Sunghyuk Park, Pavel Putrov and Nikolai Saveliev for helpful conversations and Zuzana Urbanová for the help with computer experiments. S. Gukov was partially supported by NSF grant DMS-1664227, DOE grant DE-SC0011632, and the Simons Grant New structures in low-dimensional topology. L. Katzarkov was supported by the National Science Fund of Bulgaria, National Scientific Program “VIHREN”, Project no. KP-06-DV-7 and by the Basic Research Program of the National Research University Higher School of Economics.
2. Plumbed manifolds
@@
Recall the theory of plumbed 3-manifolds. Let $\Gamma$ be a finite tree (a graph with no cycles). We call leaf a vertex of degree one and node a vertex of degree 3 and more. Additionally, we assign an integer label (framing) $m_{v}$ to each vertex $v$. Associated with this data, there is a plumbed manifold $Y$: For each vertex $v$, we take a circle bundle over 2-sphere with Euler number $m_{v}$ and then we glue along tori corresponding with the edges. One can think of $Y$ as the boundary of a 4-manifold $X$ constructed by gluing disc bundles in a similar fashion. It is convenient to describe $\Gamma$ by the plumbing matrix $M$:
$M_{vw}=\begin{cases}m_{v}&\text{if}\quad v=w\\ 1&\text{if}\quad(v,w)\in\text{Edges}(\Gamma)\\ 0&\text{otherwise.}\end{cases}$
Two different plumbing graphs represent the same 3-manifold $Y$ if and only if they are related to each other by Kirby–Neumann moves [50]; see Fig. 1.
Figure 1. Moves on plumbing trees that preserve the $3$-manifold.
The main source of plumbed manifolds are the isolated normal singularities of complex surfaces. To any such singularity germ $(X,o)$, we can assign a 3-manifold (“the link of the singularity”) by embedding a neighborhood of the singular point into $(\mathbb{C}^{n},0)$ and intersecting with a sufficiently small real sphere around zero (see e.g. [47] for details of the construction). Such plumbing graphs are always negative definite in the sense that the plumbing matrix is. Conversely, every negative definite plumbing graph is a resolution graph of some singularity and the corresponding 3-manifold is its link.
@@
The singularity $E_{12}:x^{2}+y^{3}+z^{7}=0$ has a resolution with the intersection matrix
$\left(\begin{array}[]{rrrr}-7&0&0&1\\ 0&-3&0&1\\ 0&0&-2&1\\ 1&1&1&-1\end{array}\right)$
@@
2.1. Lattices and $spin^{c}$ structures
Before we get into the $\widehat{Z}_{\sigma}(q)$ invariants, we need some preparation on lattices associated with plumbed manifold $Y$ and $spin^{c}$ structures on $Y$. We will use $\sigma,\sigma^{\prime}$ to denote $spin^{c}$ structures and $g,h,\dots$ for elements of $H=H_{1}(Y,\mathbb{Z})$. Denote $\hat{H}=Hom(H,\mathbb{C}^{*})$ the Pontryagin dual of $H.$ For a rational homology sphere $Y$, $H$ is finite and acts freely and transitively on the set of $spin^{c}$ structures $spin^{c}(Y)$, so $|spin^{c}(Y)|=|H|$. We denote the action of $h$ on $\sigma$ by $h\sigma.$
Assume now $\tilde{X}$ is a plumbed 4-manifold with boundary $Y$ and plumbing matrix $M$ of size $s\times s$. We consider the lattice $L=H_{2}(\tilde{X},\mathbb{Z})$ equipped with the intersection form. This form induces an embedding of $L$ to $L^{\prime}=H^{2}(\tilde{X},\mathbb{Z})$ and extends naturally to $L^{\prime}$. We identify the lattice $L^{\prime}$ with $(\mathbb{Z}^{s},(\vec{l},M^{-1}\vec{l}^{\prime}))$ and $L$ with $M\mathbb{Z}^{s}\subset L^{\prime}$  [^3] The basis $E^{*}_{i},i=1,\dots s$ (resp. $E_{i}$) of $L^{\prime}$ (resp. $L$) used in [47] corresponds to the vectors $-e_{i}$ (resp. $Me_{i}$) where $e_{i}$ is the standard basis of $\mathbb{Z}^{n}$. The group $H$ is then naturally identified with $L^{\prime}/L=\mathbb{Z}^{s}/M\mathbb{Z}^{s}$ using the short exact sequence:
$0\rightarrow H^{2}(\tilde{X},Y,\mathbb{Z})\rightarrow H^{2}(\tilde{X},\mathbb{Z})\rightarrow H^{2}(Y,\mathbb{Z})\rightarrow 0$
@@ 2.1. Lattices and $spin^{c}$ structures
Let $\vec{m}$ and $\vec{\delta}$ be the vectors of the framings and the degrees of the vertices, respectively. The $spin^{c}$ structures on $Y$ are naturally identified with the characteristic vectors, which are the elements of the set
$\operatorname{Char}(Y)=(2\mathbb{Z}^{s}+\vec{m})/2M\mathbb{Z}^{s}\subset L^{\prime}.$
As in [30], we use the identity $\vec{\delta}+\vec{m}=M\vec{u},$ where $\vec{u}=(1,1,\dots,1)$, to further identify $\operatorname{Char}(Y)$ with a slightly different set $\operatorname{Char}^{\prime}(Y)$:
(2.1) $\operatorname{Char}^{\prime}(Y)=(2\mathbb{Z}^{s}+\vec{\delta})/2M\mathbb{Z}^{s}\subset L^{\prime},$
via the map $\vec{l}\mapsto\vec{l}-M\vec{u}$.
The set of $spin^{c}$ structures has natural involution $\sigma\rightarrow\bar{\sigma}$ which on characteristic vectors (either $\operatorname{Char}$ or $\operatorname{Char}^{\prime}$) acts simply by $\vec{\sigma}\mapsto-\vec{\sigma}$. Fixed points of this involution, i.e., the intersection $(2\mathbb{Z}^{s}+\delta)\cap M\mathbb{Z}^{s}$, are naturally identified with $spin$ structures.
In the context of links of singularities, a special choice of $spin^{c}$ structure plays an important role [48]. This “canonical” $spin^{c}$ structure $\sigma^{\text{can}}$ has characteristic vectors $\vec{m}+2\vec{u}\in\operatorname{Char}$ and $2\vec{u}-\vec{\delta}$ in $\operatorname{Char}^{\prime}.$ A more suitable choice for our purposes is the “anticanonical” $spin^{c}$ structure $\sigma^{\text{ac}}:=\overline{\sigma^{\text{can}}{}},$ with characteristic vector $\vec{\delta}-2\vec{u}$ in $\operatorname{Char}^{\prime}.$
The choice of $spin^{c}$ structure $\sigma^{\text{ac}}$ gives us a $H$-equivariant map $H\rightarrow spin^{c}(Y)$ given by $h\rightarrow h\sigma^{\text{ac}}$. In terms of a vector representant $\vec{h}$ of $h$ in $H\cong\mathbb{Z}^{s}/M\mathbb{Z}^{s}$ and $\vec{\delta}-2\vec{u}\in\operatorname{Char}^{\prime}(Y)$, the action reads $\vec{h}\mapsto\vec{\delta}-2\vec{u}+2\vec{h}$. Note that this map preserves the involutions if and only if $\sigma^{\text{ac}}$ is a $spin$ structure. This is not the case in general.
2.2. The $q$-series invariants
@@
The $q$-series $\widehat{Z}_{\sigma}(q)$ are defined as: [^5] Let us note that by the Weyl $\mathbb{Z}_{2}$ symmetry of the integrand, we have $\widehat{Z}_{\sigma}=\widehat{Z}_{\bar{\sigma}}$ and they are often identified. We do not do it here, i.e., use the so-called “unfolded” version [8].
(2.2) $\widehat{Z}_{\sigma}(q)=q^{\frac{-3s-\text{Tr}(M)}{4}}\cdot\text{v.p.}\oint\limits_{|z_{v}|=1}\prod_{v\in\operatorname{Vert}}\frac{dz_{v}}{2\pi iz_{v}}\left(z_{v}-\frac{1}{z_{v}}\right)^{2-\deg(v)}\cdot\Theta_{\sigma}^{-M}(q,\vec{z})$
@@
(2.3) $\Theta_{\sigma}^{-M}(q,\vec{z})=\sum_{\vec{l}\in 2M\mathbb{Z}^{s}+\vec{\sigma}}q^{-\frac{(\vec{l},M^{-1}\vec{l})}{4}}\prod_{v\in\operatorname{Vert}}z_{v}^{l_{v}}.$
@@
Here, v.p. denotes taking the principal value of the integral, given by the average of the integrals over the circles $|z_{v}|=1+\epsilon$ and $|z_{v}|=1-\epsilon$, for $\epsilon>0$ small. The negativity condition is needed for the convergence of this $q$-series. The label $\sigma$ stands for a choice of $spin^{c}$ structure on $Y$, identified with a vector in $\operatorname{Char}^{\prime}(Y)=(2\mathbb{Z}^{s}+\vec{\delta})/2M\mathbb{Z}^{s}$; see (2.1).
@@
We can reformulate the integral formula using symmetric expansions. Recall that for a rational function $r(t)\in\mathbb{C}(t)$, the symmetric expansion of $r(t)$, denoted $\operatorname{s.e.}r(t)$, is the average of Laurent expansions around 00 and $\infty$:
@@
$\operatorname{s.e.}r(t):=\frac{1}{2}\left(\operatorname*{\text{expn\,}}_{t\rightarrow 0}r(t)+\operatorname*{\text{expn\,}}_{t\rightarrow\infty}r(t)\right)$
Using this notion, we can turn (2.2) into the following formula
(2.4) $\widehat{Z}_{\sigma}(q)=2^{-s}q^{\frac{-3s-\text{Tr}(M)}{4}}\mathcal{L}_{-M}(F_{\sigma}(\vec{z}\,)),$
where $F_{\sigma}(\vec{z}\,)=F_{\sigma}(z_{1},\dots,z_{s})\in\mathbb{Z}[[z_{1}^{\pm 1},\dots,z_{s}^{\pm 1}]]$ are formal power series
@@
(2.5) $F_{\sigma}(\vec{z}\,)=\pi_{\sigma}\left(\prod_{v\,\in\,\operatorname{Vert}}2\operatorname{s.e.}\left(z_{v}-\frac{1}{z_{v}}\right)^{2-\delta_{v}}\right),$
where $\pi_{\sigma}$ is the projection on the class of $\sigma$:
$\pi_{\sigma}(\vec{z}\,^{\vec{l}})=\begin{cases}\vec{z}\,^{\vec{l}}&\text{if }[\vec{l}\,]=\sigma\\ 0&\text{otherwise}.\end{cases}$
@@
$\mathcal{L}_{-M}$ denotes the following transformation (“Laplace transform”)
(2.6) $\mathcal{L}_{-M}:\vec{z}^{\,\vec{l}}\rightarrow q^{\frac{-(\vec{l},M^{-1}\vec{l})}{4}},$
extended linearly to formal power series in $\vec{z}^{\pm 1}.$
We can eliminate the projections using the $spin^{c}$ structure $\sigma^{\text{ac}}$:
@@
(2.7) $\sum_{h\in H}F_{h\sigma^{\text{ac}}}(\vec{z}\,)h=\prod_{v\,\in\,\operatorname{Vert}}z_{i}^{\delta_{v}-2}\prod_{v\,\in\,\operatorname{Vert}}2\operatorname{s.e.}(g_{v}z_{v}^{2}-1)^{2-\delta_{v}},$
@@
where $g_{v}$ are the generators of $H$ given by classes $[e_{v}]$ of canonical basis vectors of $\mathbb{Z}^{s}/M\mathbb{Z}^{s}.$ Both sides of this equation should be understood as elements of the group ring $\mathbb{Z}[\vec{z}^{\,\pm 1}\,][H]$ of $H,$ with coefficients in formal power series in $\vec{z}$ and the symmetric expansion is naturally extended to $\mathbb{C}(t)[H]$.
3. $Z_{0}$ and splice diagrams

### v3

@@ 1.2 Spectra in algebra and geometry
Since one of the predictions of the so-called 3d-3d correspondence is that a closed 3-manifold $Y$ corresponds to a VOA [8, 9, 68] for which $\widehat{Z}_{\sigma}$ is a character, it is natural to ask how the spectrum of conformal weights $\Delta_{\sigma}$ relates to other spectra mentioned above, defined more geometrically. Contrary to what one might naively expect, we find that the spectrum of conformal weights $\Delta_{\sigma}$ is rather different from the spectrum of the corresponding hypersurface singularity: while the former determines the leading $q$-power in the $q$-expansion, the latter encodes the structure of the $q$-series coefficients at large $q$-powers. In turn, the latter determines the behavior of the $q$-series near $q={\rm e}^{\hbar}\approx 1$ or, equivalently, the expansion in $\hbar$ near $\hbar=0$.
To summarize, expansion in $q$, as in $\widehat{Z}_{\sigma}(q)$, has direct contact with counting problems (curve counting, BPS states, etc.), with vertex algebras, and with the spectrum of conformal dimensions. These connections become less natural near $q={\rm e}^{\hbar}\approx 1$, where connections to complex Chern–Simons theory and the spectrum of hypersurface singularities become manifest. The two expansion limits are related by resurgent analysis, a powerful technique that, roughly speaking, allows to transfer (enlarge) the domain of a given function (or a power series, possibly, with zero radius of convergence).
This interplay between expansions near $q=0$ and $q={\rm e}^{\hbar}\approx 1$ plays an important role in curve counting [17, 28, 37, 42] and — since $\widehat{Z}_{\sigma}(q)$ admit an interpretation via curve counting too [20, 21, 33, 34] — in the study of $\widehat{Z}_{\sigma}(q)$-invariants [3, 11, 13, 30, 70]. In particular, near $q={\rm e}^{\hbar}\approx 1$ more natural objects are $Z_{a}(q)$, without a ‘hat’, and trans-series $Z_{\alpha}(q)$ of the complex Chern–Simons theory. This is consistent with what we find here: $Z_{0}$ and, more generally, $Z_{\alpha}$ appear to be more natural from the perspective of the singularity theory. And, perhaps not surprisingly (given the above explanations), the spectrum of the hypersurface singularity associated with $Y$ is closely related to the set of values of the classical Chern–Simons functional on $Y$, i.e., to the singularities on the Borel plane.
2 Plumbed manifolds
@@
Recall the theory of plumbed 3-manifolds. Let $\Gamma$ be a finite tree (a graph with no cycles). We call leaf a vertex of degree one and node a vertex of degree 3 and more. Additionally, we assign an integer label (framing) $m_{v}$ to each vertex $v$. Associated with this data, there is a plumbed manifold $Y$: For each vertex $v$, we take a circle bundle over 2-sphere with Euler number $m_{v}$ and then we glue along tori corresponding with the edges. One can think of $Y$ as the boundary of a 4-manifold $X$ constructed by gluing disc bundles in a similar fashion. It is convenient to describe $\Gamma$ by the plumbing matrix $M$
$M_{vw}=\begin{cases}m_{v}&\text{if}\ v=w,\\ 1&\text{if}\ (v,w)\in\text{Edges}(\Gamma),\\ 0&\text{otherwise}.\end{cases}$
Two different plumbing graphs represent the same 3-manifold $Y$ if and only if they are related to each other by Kirby–Neumann moves [50]; see Figure 1.
Figure 1: Moves on plumbing trees that preserve the 33-manifold.
The main source of plumbed manifolds are the isolated normal singularities of complex surfaces. To any such singularity germ $(X,o)$, we can assign a 3-manifold (“the link of the singularity”) by embedding a neighborhood of the singular point into $(\mathbb{C}^{n},0)$ and intersecting with a sufficiently small real sphere around zero (see, e.g., [47] for details of the construction). Such plumbing graphs are always negative definite in the sense that the plumbing matrix is. Conversely, every negative definite plumbing graph is a resolution graph of some singularity and the corresponding 3-manifold is its link.
@@
The singularity $E_{12}\colon x^{2}+y^{3}+z^{7}=0$ has a resolution with the intersection matrix
$\begin{pmatrix}-7&0&0&1\\ 0&-3&0&1\\ 0&0&-2&1\\ 1&1&1&-1\end{pmatrix}$.
@@
2.1 Lattices and $\boldsymbol{{\rm spin}^{c}{}}$ structures
Before we get into the $\widehat{Z}_{\sigma}(q)$ invariants, we need some preparation on lattices associated with plumbed manifold $Y$ and ${\rm spin}^{c}$ structures on $Y$. We will use $\sigma$, $\sigma^{\prime}$ to denote ${\rm spin}^{c}$ structures and $g,h,\dots$ for elements of $H=H_{1}(Y,\mathbb{Z})$. Denote $\hat{H}=\operatorname{Hom}(H,\mathbb{C}^{*})$ the Pontryagin dual of $H$. For a rational homology sphere $Y$, $H$ is finite and acts freely and transitively on the set of ${\rm spin}^{c}$ structures ${\rm spin}^{c}(Y)$, so $|{\rm spin}^{c}(Y)|=|H|$. We denote the action of $h$ on $\sigma$ by $h\sigma$.
[…]
$0\rightarrow H^{2}\bigl{(}\tilde{X},Y,\mathbb{Z}\bigr{)}\rightarrow H^{2}\bigl{(}\tilde{X},\mathbb{Z}\bigr{)}\rightarrow H^{2}(Y,\mathbb{Z})\rightarrow 0$
@@ 2.1 Lattices and $\boldsymbol{{\rm spin}^{c}{}}$ structures
Let $\vec{m}$ and $\vec{\delta}$ be the vectors of the framings and the degrees of the vertices, respectively. The ${\rm spin}^{c}$ structures on $Y$ are naturally identified with the characteristic vectors, which are the elements of the set $\operatorname{Char}(Y)=(2\mathbb{Z}^{s}+\vec{m})/2M\mathbb{Z}^{s}$. As in [29], we use the identity $\vec{\delta}+\vec{m}=M\vec{u},$ where $\vec{u}=(1,1,\dots,1)$, to further identify $\operatorname{Char}(Y)$ with a slightly different set $\operatorname{Char}^{\prime}(Y)$
$\operatorname{Char}^{\prime}(Y)=\bigl{(}2\mathbb{Z}^{s}+\vec{\delta}\bigr{)}/2M\mathbb{Z}^{s},$ (2.1)
via the map induced by $\vec{l}\mapsto\vec{l}-M\vec{u}$.
The set of ${\rm spin}^{c}$ structures has natural involution $\sigma\rightarrow\bar{\sigma}$ which on characteristic vectors (either $\operatorname{Char}$ or $\operatorname{Char}^{\prime}$) acts simply by $\vec{\sigma}\mapsto-\vec{\sigma}$. Fixed points of this involution, i.e., the intersection $(2\mathbb{Z}^{s}+\delta)\cap M\mathbb{Z}^{s}$}, are naturally identified with spin structures.
In the context of links of singularities, a special choice of ${\rm spin}^{c}$ structure plays an important role [48]. This “canonical” ${\rm spin}^{c}$ structure $\sigma^{{\rm can}}$ has characteristic vectors $\vec{m}+2\vec{u}\in\operatorname{Char}$ and $2\vec{u}-\vec{\delta}$} in $\operatorname{Char}^{\prime}$. A more suitable choice for our purposes is the “anticanonical” ${\rm spin}^{c}$ structure $\sigma^{{\rm ac}}:=\overline{\sigma^{{\rm can}}{}},$ with characteristic vector $\vec{\delta}-2\vec{u}$ in $\operatorname{Char}^{\prime}$.
The choice of ${\rm spin}^{c}$ structure $\sigma^{{\rm ac}}$ gives us a $H$-equivariant map $H\rightarrow{\rm spin}^{c}(Y)$ given by $h\rightarrow h\sigma^{{\rm ac}}$. In terms of a vector representative $\vec{h}$ of $h$ in $H\cong\mathbb{Z}^{s}/M\mathbb{Z}^{s}$ and $\vec{\delta}-2\vec{u}\in\operatorname{Char}^{\prime}(Y)$, the action reads $\vec{h}\mapsto\vec{\delta}-2\vec{u}+2\vec{h}$. Note that this map preserves the involutions if and only if $\sigma^{{\rm ac}}$ is a spin structure. This is not the case in general.
2.2 The $\boldsymbol{q}$-series invariants
@@
The $q$-series $\widehat{Z}_{\sigma}(q)$ are defined as [^7] Let us note that by the Weyl $\mathbb{Z}_{2}$ symmetry of the integrand, we have $\widehat{Z}_{\sigma}=\widehat{Z}_{\bar{\sigma}}$ and they are often identified. We do not do it here, i.e., use the so-called “unfolded” version [8].
$\widehat{Z}_{\sigma}(q)=q^{\frac{-3s-\text{Tr}(M)}{4}}\cdot\text{v.p.}\oint\limits_{|z_{v}|=1}\prod_{v\in\operatorname{Vert}}\frac{{\rm d}z_{v}}{2\pi{\rm i}z_{v}}\left(z_{v}-\frac{1}{z_{v}}\right)^{2-\deg(v)}\cdot\Theta_{\sigma}^{-M}(q,\vec{z}),$ (2.2)
@@
$\Theta_{\sigma}^{-M}(q,\vec{z})=\sum_{\vec{l}\in 2M\mathbb{Z}^{s}+\vec{\sigma}}q^{-\frac{(\vec{l},M^{-1}\vec{l})}{4}}\prod_{v\in\operatorname{Vert}}z_{v}^{l_{v}}.$
@@
Here, v.p. denotes taking the principal value of the integral, given by the average of the integrals over the circles $|z_{v}|=1+\epsilon$ and $|z_{v}|=1-\epsilon$, for $\epsilon>0$ small. The negativity condition is needed for the convergence of this $q$-series. The label $\sigma$ stands for a choice of ${\rm spin}^{c}$ structure on $Y$, identified with a vector in $\operatorname{Char}^{\prime}(Y)=\bigl{(}2\mathbb{Z}^{s}+\vec{\delta}\,\bigr{)}/2M\mathbb{Z}^{s}$; see (2.1).
@@
We can reformulate the integral formula using symmetric expansions. Recall that for a rational function $r(t)\in\mathbb{C}(t)$, the symmetric expansion of $r(t)$, denoted $\mathrm{s.e.}\,r(t)$, is the average of Laurent expansions around 0 and $\infty$
@@
$\mathrm{s.e.}\,r(t):=\frac{1}{2}\bigl{(}\operatorname*{\text{expn\,}}_{t\rightarrow 0}r(t)+\operatorname*{\text{expn\,}}_{t\rightarrow\infty}r(t)\bigr{)}.$
Using this notion, we can turn (2.2) into the following formula:
$\widehat{Z}_{\sigma}(q)=2^{-s}q^{\frac{-3s-\text{Tr}(M)}{4}}\mathcal{L}_{-M}(F_{\sigma}(\vec{z})),$
where $F_{\sigma}(\vec{z})=F_{\sigma}(z_{1},\dots,z_{s})\in\mathbb{Z}\big{[}\big{[}z_{1}^{\pm 1},\dots,z_{s}^{\pm 1}\big{]}\big{]}$ are formal power series
@@
$F_{\sigma}(\vec{z})=\pi_{\sigma}\left(\prod_{v\in\operatorname{Vert}}2\mathrm{s.e.}\,\left(z_{v}-\frac{1}{z_{v}}\right)^{2-\delta_{v}}\right),$ (2.3)
where $\pi_{\sigma}$ is the projection on the class of $\sigma$
$\pi_{\sigma}\bigl{(}\vec{z}\,{}^{\vec{l}}\bigr{)}=\begin{cases}\vec{z}\,{}^{\vec{l}}&\text{if }[\vec{l}]=\sigma,\\ 0&\text{otherwise}.\end{cases}$
@@
$\mathcal{L}_{-M}$ denotes the following transformation (“Laplace transform”) $\mathcal{L}_{-M}\colon\vec{z}\,{}^{\vec{l}}\rightarrow q^{\frac{-(\vec{l},M^{-1}\vec{l})}{4}}$, extended linearly to formal power series in $\vec{z}\,{}^{\pm 1}$.
We can eliminate the projections using the ${\rm spin}^{c}$ structure $\sigma^{{\rm ac}}$
@@
$\sum_{h\in H}F_{h\sigma^{{\rm ac}}}(\vec{z})h=\prod_{v\in\operatorname{Vert}}z_{v}^{\delta_{v}-2}\prod_{v\in\operatorname{Vert}}2\mathrm{s.e.}\,\bigl{(}g_{v}z_{v}^{2}-1\bigr{)}^{2-\delta_{v}},$ (2.4)
@@
where $g_{v}$ are the generators of $H$ given by classes $e_{v}]$ of canonical basis vectors of $\mathbb{Z}^{s}/M\mathbb{Z}^{s}$. Both sides of this equation should be understood as elements of the group ring $\mathbb{Z}[\vec{z}\,{}^{\pm 1}][H]$ of $H,$ with coefficients in formal power series in $\vec{z}$ and the symmetric expansion is naturally extended to $\mathbb{C}(t)[H]$.
3 $\boldsymbol{Z_{0}}$ and splice diagrams